OpenStudy (anonymous):

Find the critical points of f(x) = x^4*(x-1)^3

OpenStudy (anonymous):

you first need to find the derivative of the equation then set the equation to zero solve for x and you will have your critical points

OpenStudy (anonymous):

\[f(x) =x^{4}(x-1)^{3} \] is this what your equation looks like?

OpenStudy (anonymous):

Yes, it is

OpenStudy (anonymous):

I know to do the product rule, but I get stuck after that.

OpenStudy (anonymous):

Why did you use the product rule? I was thinking chain rule ? hmm

OpenStudy (anonymous):

I end up with: \[x ^{4}3(x-1)^{2}+4x ^{3}(x-1)^{3}\]

OpenStudy (anonymous):

Chain rule confuses me quickly. I can understand it with y' notation, but I still can't grasp dy/dx and all of that. I have less than a month left and no more Calculus classes after this.

OpenStudy (anonymous):

dy/dx is the same thing as y' (y prime)

OpenStudy (anonymous):

I know, but I get lost quickly with the dy over dx and such. Never understood it. I've gotten by with y' notation. Looks cleaner to me

OpenStudy (anonymous):

chain rule is simply f'(g(x)) * (g'(x)) so basically f'=derivative of what's outside multiplied by whats inside (leave the inside alone) and then take derivative of inside... does that help?

OpenStudy (anonymous):

yes y' is better to most folks.

OpenStudy (anonymous):

I do get it to an extent. I get lost as to how it applies to this problem.

OpenStudy (perl):

i like u substitution myself

OpenStudy (perl):

example u^n -> n*u^(n-1) * u' and u ' = du/dx

OpenStudy (perl):

u substitution captures the 'form' of it

OpenStudy (anonymous):

So \[(x-1)^{3}\] becomes \[3(x-1)^{2}* 1\]

OpenStudy (perl):

3(x-1)^2 * (x-1)' 3(x-1)^2 * 1

OpenStudy (anonymous):

Okay, but what happens with the \[x ^{4}\] here?

OpenStudy (perl):

there you have 4u^3 * u' , so here u = x 4(x)^3 * x' 4x^3 * 1

OpenStudy (perl):

so u substitution works with just x, but you can use the simple power rule when it is just x

OpenStudy (anonymous):

Okay

OpenStudy (perl):

i use the u rule when i have something more complicated than a function with x alone

OpenStudy (perl):

(x^2 - 3x) ^5 , the derivative of that is hard using only simple power rule. i would have to expand that

OpenStudy (anonymous):

I understand. So for my original problem, I still would use the product rule with the chain rule, right?

OpenStudy (nincompoop):

not the right wording

OpenStudy (nincompoop):

let us try to recall what the general product rule is for simplicity I will use the u'v + uv' which is the same when we split it into two functions f(x) and g(x) f'(x) * g(x) + f(x) * g'(x) so we are looking at two derivatives - one for the u and one for the v in your problem, f(x) = x^4(x-1)^3, in order to get the derivative, you need to identify your u and your v u = x^4 v = (x-1)^3 now, we obtain their derivatives individually: u requires only a power-rule 4x^3 v requires the chain-rule 3(x-1)^2 * 1 = 3(x-1)^2 if you're getting confused with chain-rule, think of it as the big fish and the small fish. big fish being the outermost of the expression without touching the contents inside the parenthesis and the small fish is whatever you have inside the parenthesis. so we obtained the big fish's derivative first then multiply it by the small fish's derivative. so, we clean things up so you can see what we have so far PRODUCT RULE u'v + uv' u = x^4, u' = 4x^3 v = (x-1)^3, v' = 3(x-1)^2 now apply our product rule